OpenStudy (anonymous):

A nurse gave a patient gr xiiss of Aminophylline. Tablets available are 250 mg. How many tablets should the nurse give the patient?

OpenStudy (anonymous):

What does gr xiiss mean?

OpenStudy (anonymous):

gr= gram ss = 1/2 xii I am puzzled

OpenStudy (welshfella):

looks a bit like roman numerals Though is s a roman numeral? xii in roman numerals is 12

OpenStudy (anonymous):

that's what I thought too

OpenStudy (welshfella):

- though it might be something completely different

OpenStudy (anonymous):

grams 12 1/2 i guess

OpenStudy (welshfella):

thats is 12,500 mg / 250 is a lot of tablets!!

OpenStudy (acxbox22):

overdose

OpenStudy (anonymous):

that is what I thought too, in reality that will be a lot to swallow

OpenStudy (welshfella):

could gr mean grain?

OpenStudy (anonymous):

...exactrly

OpenStudy (acxbox22):

is that the exact question as stated?

OpenStudy (welshfella):

25!

OpenStudy (anonymous):

gr could be grain too

OpenStudy (anonymous):

yes it is the exact question

OpenStudy (texaschic101):

xiiss = 12.5....there are 60 mg I one grain...so grxiis = 750 mg

OpenStudy (anonymous):

i think grain will be realistic right

OpenStudy (welshfella):

no 50 if its 12 1/2 grams

OpenStudy (welshfella):

yea graons would be right , i think

OpenStudy (anonymous):

750/250=?

OpenStudy (texaschic101):

I think so ^^

OpenStudy (welshfella):

thats defnitely more realistic!

OpenStudy (anonymous):

@ninamae05 want to finish it?

OpenStudy (anonymous):

1 grain = 60 mg 12 1/2 grains x 60 mg / 1 grain = 750 mg 250 mg = 750 mg --------- --------- 1 tab x 250 x = 750 x = 3 tabs

OpenStudy (anonymous):

Seems to be true to me.

OpenStudy (welshfella):

yea i would go with that.

OpenStudy (texaschic101):

me too ^

OpenStudy (anonymous):

http://www.mgccc.edu/learning_lab/science/nursing/level1.html This might also be helpful in the future. ^

OpenStudy (welshfella):

learnt something new 1 grain = 60 mg